# A Survey on Cross-Domain Sequential Recommendation

### v2

@@ Abstract
Cross-domain sequential recommendation (CDSR) shifts the modeling of user preferences from flat to stereoscopic by integrating and learning interaction information from multiple domains at different granularities (ranging from inter-sequence to intra-sequence and from single-domain to cross-domain). In this survey, we first define the CDSR problem using a four-dimensional tensor and then analyze its multi-type input representations under multidirectional dimensionality reductions. Following that, we provide a systematic overview from both macro and micro views. From a macro view, we abstract the multi-level fusion structures of various models across domains and discuss their bridges for fusion. From a micro view, focusing on the existing models, we specifically discuss the basic technologies and then explain the auxiliary learning technologies. Finally, we exhibit the available public datasets and the representative experimental results as well as provide some insights into future directions for research in CDSR.
@@ 1 Introduction
In real life, people leave interaction traces in multiple scenarios (e.g., shopping scenarios, reading scenarios, etc.), multiple platforms (e.g., Amazon, Taobao, etc.), and even multiple boards (e.g., books and movies in Douban, etc.) all the time. Considering all these as different domains, it is a critical step for the development of recommender systems to combine information across domains so that the data of user preferences is no longer fragmented, and shifting the modeling from flat to stereoscopic. As shown in Table 1, recommender systems focus on information that is becoming deeper and more comprehensive, starting from traditional one-class collaborative filtering (OCCF), further progressing to sequential one-class collaborative filtering (SOCCF) and to cross-domain one-class collaborative filtering (CD-OCCF), finally to cross-domain sequential one-class collaborative filtering (CD-SOCCF, also called cross-domain sequential recommendation, CDSR).
@@
In the case of OCCF Rendle et al. (2012), the available information is limited to whether the user interacted with the item. As illustrated in Figure 1, we only know all the items that user 1 and user 2 have interacted with in a single domain, without considering the sequential pattern among successive items. SOCCF Lin et al. (2020), building upon OCCF, places greater emphasis on the order relationships between items within a sequence, i.e., the information of intra-sequence in a single domain, and thus the model can capture the user’s long-term and short-term preferences. At a large granularity, expanding from a single domain to multiple domains offers a fresh perspective to address sparse data issues in recommender systems, thus constituting CD-OCCF Hu et al. (2018). For instance, as shown in Figure 1, users have interactions not only in the book domain but also in the closely related movie domain. In the cross-domain scenario, we must meet the challenges that transferring or aggregating interaction records from multiple domains as well as aligning representations of users or items across different domains.
Indeed, CDSR combines all of the aforementioned directions and further incorporates sequential information based on CD-OCCF and cross-domain information based on SOCCF. As indicated by the purple arrows in Figure 1, CDSR necessitates the precise capturing of sequential relationships between items across different domains. For example, user 1, after reading two books in the $Harry\ Potter$ series, chooses to watch the corresponding $Harry\ Potter$ movie. Similarly, user 2 who enjoys watching romantic movies also reads the original novels after watching the $Pride\ and\ Prejudice$ movie. Besides, it is also worth considering in CDSR how to fuse sequential information from different domains as well as how to distinguish between users’ specific preferences within a single domain and global preferences shared across multiple domains.
In this survey, we first formulate CDSR’s problem definition and modeling, considering various dimensionality reductions and different input representations. Then we adopt both a macro and micro view to summarize the existing works in CDSR. From a macro view, we present an overview of multi-level fusion structures, discussing how to fuse information across different domains and exploring bridges for cross-domain fusion. From a micro view, we conduct a detailed analysis of various technologies employed by existing works that are categorized into basic technologies and auxiliary learning. Based on this analysis, we synthesize a comprehensive classification table (as shown in Table 2) and framework figure of the key technologies (as shown in Figure 5). Furthermore, we also list the datasets commonly used in CDSR and the representative experimental results as well as provide some insights into potential future directions.
@@ 2 Formulation
In this section, we first present the problem definition that unifies CDSR using a four-dimensional tensor. Then, referring to Figure 2, we delve into the various directions of dimensionality reduction that are often applied in practical modeling and the corresponding multi-type input representations.
@@ 2.1 Problem Definition
Notice that most existing works in CDSR are primarily established on two domains, i.e., $k=2$. Therefore, for the subsequent discussion, we take two domains as an example and represent them as domain A and domain B, respectively. In fact, some researchers Bi et al. (2020); Xu et al. (2023c) also refer to the two domains as the source domain and the target domain. Moreover, some works Ma et al. (2019); Sun et al. (2023); Guo et al. (2021, 2023c) introduce an additional assumption with a shared account. They assume that within a real account, there are $q$ virtual users ($u_{j},j\leq q$) simultaneously active. Regarding the shared account issue, it is applicable not only in cross-domain scenarios but also in single-domain scenarios.
@@ 2.2 Multidirectional Dimensionality Reduction
For computational and modeling convenience, the four-dimensional tensor is often reduced to three-dimensional tensors from various directions (as shown in Figure 2). Below, we explain dimensionality reduction from three directions:
  • W/o User Dimension. For each user, their interactions in chronological order can form a sequence in each domain. These sequences reflect each user’s personalization as well as long and short-term preferences in different domains, and can further serve as the basis for constructing a local graph in subsequent work.
@@
  • W/o Domain Dimension. Blending two or more domains into a hybrid domain, so that the models can clarify the passing relationships between items originally in different domains. Many existing studies treat the hybrid domain as an independent domain, designing model fusion structures that incorporate all three domains in parallel, which is shown in Section 3.1.
@@ 2.3.1 Pure Sequential Representation
Arranging the interacted items of a user in chronological order yields a sequence. For each user $u$, we define $S^{A}_{u}=\{i^{A}_{1},i^{A}_{2},\dots\}$ and $S^{B}_{u}=\{i^{B}_{1},i^{B}_{2},\dots\}$ as his/her interaction sequences in domain A and domain B, respectively. If $u$ is an overlapping user who has interactions both in domain A and domain B, we can mix $S^{A}_{u}$ and $S^{B}_{u}$ to a hybrid sequence in chronological order (e.g., $S^{hybrid}_{u}=\{i^{A}_{1},i^{B}_{1},i^{A}_{2},i^{A}_{3},i^{B}_{2},\dots\}$), which is viewed as a sequence of the hybrid domain. From a pure sequential perspective, we can substitute the four-dimensional tensor in Eq. (1) with the set of sequences in each domain, as follows,
@@ 2.3.2 Graph-Encoded Sequential Representation
Some researchers Guo et al. (2021); Zheng et al. (2022); Cao et al. (2022); Zhang et al. (2023b) turn to construct directed graphs $G=\{V,E\}$ to model sequential information, where $V$ is a set of items that have been interacted with and $E$ is the edges that represent relations of the serial relationship from item to item. Constructing the user’s sequence within a domain or the sequence within a session as a local graph is a common practice Zheng et al. (2022); Chen et al. (2021). In the global graph construction scenario, $E$ is also utilized to denote the relations between users and items Xu et al. (2023c). We represent the raw data in its graph-encoded form to extend Eq.(1), where $G_{l}$ denotes all local graphs and $G_{g}$ denotes the global graph, as follows,
@@ 2.3.3 Side Information
More and more researchers consider incorporating side information to enrich the semantic representation of users’ historical behaviors. Here, we divide the most common side information into three categories: time, text (i.e., user/item context and review), and knowledge graph.
Time. Previously the timestamps are only used to order the items, but some researchers Xiao et al. (2023); Wang et al. (2022); Guo et al. (2022) define $t_{ij}=|t_{i}-t_{j}|$ to model the time interval and apply it to the subsequent learning of model, where $t_{i}$ and $t_{j}$ are the timestamps that the item $i$ and the item $j$ are interacted with, respectively. Besides, we can also mine more information about time, such as periodicity and the duration of the interaction, etc.
@@
Knowledge Graph. Knowledge graphs Bi et al. (2020); Ma et al. (2022) deal with information from both entity and relationship views. It not only contains structural information between nodes but also implies some relationship description. A knowledge graph defines an entity set $E^{KG}$ and a relation set $R^{KG}$, which consists of multiple entity-relation-entity triples $<e_{i},r,e_{j}>$ (e.g., $<e_{i_{1}^{A}},Is\_the\_same\_category,e_{i_{1}^{B}}>$ meaning that the entity $e_{i_{1}^{A}}$ from domain A has the same category as entity $e_{i_{1}^{B}}$ from domain B).
Considering the mentioned side information, we can extend Eq. (1) as follows,
@@ 2.3.4 Pre-trained Features
Considering the privacy issues in two or more domains, some works Ding et al. (2023); Zhang et al. (2023a); Lei et al. (2021) use model-trained features as the input of the source domain, instead of raw data. Ding et al. (2023) directly combines the pre-trained user features from one domain with another domain. Zhang et al. (2023a) utilizes federated learning to fetch the global representation from the server, and then injects it into the local model.
@@ 3 Macro-View: What Structure Is Used to Fuse the CDSR Information?
In CDSR, the fusion structure serves as the primary skeleton of the model, providing pathways for the separation and aggregation of these features. In this section, we first propose the multi-level fusion structures and then elaborate on the bridge for inter-domain fusion.
@@ 3.1.1 One-Level Fusion Structures
In earlier works, to combine the cross-domain information, there are a lot of works that first learn user preferences from domain A and domain B separately, and then fuse the representations of the two domains using various operations (shown in Figure 3(a)). These operations are not limited to concatenation Lei et al. (2021); Guo et al. (2023c), summation Alharbi and Caragea (2021); Ding et al. (2023), multi-layer perceptron (MLP) Bi et al. (2020), or some attention mechanisms Ouyang et al. (2020); Li et al. (2021).
Additionally, some works employ transfer learning to transfer knowledge from a source domain to a target domain Chen et al. (2021); Liu and Zhu (2021), or train a discriminator to bridge the representation of two domains with the idea of adversarial learning Li et al. (2022), shown in Figure 3(b).
@@ 3.1.2 Two/Three-Level Fusion Structures
Considering the potential for further exploration and advancement in the field of fusion structure, some works combine the single domain and the hybrid domain with a multi-level fusion structure. As shown in Figure 3(e), in order to predict the next item in a separate domain, the hybrid domain is utilized as the main sharer Cao et al. (2022); Wang et al. (2022) or the bridge Zheng et al. (2022) to transfer knowledge from another domain. There are also some works Ye et al. (2023) that choose to combine the domain A and domain B first and then fuse the hybrid information (i.e., Figure 3(f)).
@@ 3.1.3 Discussion
Although we propose multi-level fusion structures, it does not mean that a more complex structure will result in a better performance. The simple fusion structures, i.e., one-level fusion structures, are easy to implement but may not be conducive to comprehensive modeling of domain-specific and domain-generic features due to the limitation of the degree of fusion. While improving effectiveness, multi-level fusion structures bring increased complexity and reduced interpretability.
@@ 3.2 Bridges for Cross-Domain Fusion
When comes to cross-domain fusion, it’s vital to clarify the bridges for information sharing between domains. We divide those bridges into three categories: the same users, contexts, and clusters/groups, as shown in Figure 4. The first is prevalent in scenarios with overlapping users, while the latter two apply in scenarios without overlapping users.
  • The Same Users. If the same users exist, then they are the first choice as the bridge. Some works only focus on fully overlapping users, neglecting non-overlapping users, while others Li et al. (2022) consider both cases.
  • Contexts. If the same users don’t exist, it is feasible to leverage the semantic similarity of natural language. For instance, we can replace traditional item IDs with item contexts to explore similarities between items of different domains Liu et al. (2023).
@@
Different bridges exhibit distinct characteristics, each suitable for different scenarios. Works relying on overlapping users often perform well but their application scenarios may face limitations due to sparse real-world data. Conversely, works built on non-overlapping users can be applied to a broader range of scenarios but may have limited performance.
@@ 4 Micro-View: What Technologies Are Used to Address the CDSR Problem?
Below, we take a more tangible perspective to summarize the technologies adopted by existing models in addressing the challenges. In conjunction with Table 2 and Figure 5, we elaborate on the basic technologies and the auxiliary learning technologies, respectively.
@@ 4.1.2 Attention Mechanisms in CDSR
[…]

### v3

@@ Abstract
Cross-domain sequential recommendation (CDSR) shifts the modeling of user preferences from flat to stereoscopic by integrating and learning interaction information from multiple domains at different granularities (ranging from inter-sequence to intra-sequence and from single-domain to cross-domain). In this survey, we first define the CDSR problem using a four-dimensional tensor and then analyze its multi-type input representations under multidirectional dimensionality reductions. Following that, we provide a systematic overview from both macro and micro views. From a macro view, we abstract the multi-level fusion structures of various models across domains and discuss their bridges for fusion. From a micro view, focusing on the existing models, we first discuss the basic technologies and then explain the auxiliary learning technologies. Finally, we exhibit the available public datasets and the representative experimental results as well as provide some insights into future directions for research in CDSR.
@@ 1 Introduction
In real life, people leave interaction traces in multiple scenarios (e.g., shopping scenarios, reading scenarios, etc.), multiple platforms (e.g., Amazon, Taobao, etc.), and even multiple boards (e.g., books and movies in Douban, etc.) all the time. Considering all these as different domains, it is a critical step for developing recommender systems to combine information across domains so that the data of user preferences is no longer fragmented, shifting the modeling from flat to stereoscopic. As shown in Table 1, recommender systems focus on information that is becoming deeper and more comprehensive, starting from traditional one-class collaborative filtering (OCCF), further progressing to sequential one-class collaborative filtering (SOCCF) and cross-domain one-class collaborative filtering (CD-OCCF), and finally to cross-domain sequential one-class collaborative filtering (CD-SOCCF, also called cross-domain sequential recommendation, CDSR).
@@
In the case of OCCF Rendle et al. (2012), the available information is limited to whether a user interacts with an item. As illustrated in Figure 1, we only know all the items that user 1 and user 2 have interacted with in a single domain, without considering the sequential pattern among successive items. SOCCF Lin et al. (2020), building upon OCCF, places greater emphasis on the order relationships between items within a sequence, i.e., the information of intra-sequence in a single domain, and thus the model can capture the user’s long-term and short-term preferences. At a large granularity, expanding from a single domain to multiple domains offers a fresh perspective to address the data sparsity issue in recommender systems, thus constituting CD-OCCF Hu et al. (2018). For instance, as shown in Figure 1, users have interactions not only in the book domain but also in the closely related movie domain. In a cross-domain scenario, we must meet the challenges of transferring or aggregating interaction records from multiple domains and aligning representations of users or items across different domains.
Indeed, CDSR combines all of the aforementioned directions and further incorporates sequential information based on CD-OCCF and cross-domain information based on SOCCF. As indicated by the purple arrows in Figure 1, CDSR necessitates the precise capturing of sequential relationships between items across different domains. For example, user 1, after reading two books in the $Harry\ Potter$ series, chooses to watch the corresponding $Harry\ Potter$ movie. Similarly, user 2, who enjoys watching romantic movies, also reads the original novels after watching the $Pride\ and\ Prejudice$ movie. Besides, it is also worth considering in CDSR how to fuse sequential information from different domains as well as how to distinguish between users’ specific preferences within a single domain and global preferences shared across multiple domains.
In this survey, we first formulate the CDSR problem and modeling tasks, considering various dimensionality reductions and different input representations. Then we adopt both macro and micro views to summarize the existing works in CDSR. From a macro view, we present an overview of multi-level fusion structures, discussing how to fuse information across different domains and exploring bridges for cross-domain fusion. From a micro view, we conduct a detailed analysis of various technologies employed by existing works that are categorized into basic and auxiliary learning technologies. Based on this analysis, we obtain a comprehensive classification shown in Table 2 and framework figure of the key technologies ( shown in Figure 5). Furthermore, we list the datasets commonly used in CDSR and the representative experimental results as well as provide some insights into potential future directions.
@@ 2 Formulation
In this section, we first present the problem definition of CDSR using a four-dimensional tensor. Then, referring to Figure 2, we delve into the various directions of dimensionality reduction that are often adopted in practical modeling and the corresponding multi-type input representations.
@@ 2.1 Problem Definition
Notice that most existing works in CDSR are primarily established on two domains, i.e., $k=2$. Therefore, for the subsequent discussion, we take two domains as an example and represent them as domain A and domain B, respectively. In fact, some researchers Bi et al. (2020); Xu et al. (2023c) also refer to the two domains as a source domain and a target domain. Moreover, some works Ma et al. (2019); Sun et al. (2023); Guo et al. (2021, 2023c) introduce an additional assumption with a shared account. They assume that within a real account, there are $q$ virtual users ($u_{j},j\leq q$) simultaneously active. Regarding the shared account issue, it is applicable not only in cross-domain scenarios but also in single-domain scenarios.
@@ 2.2 Multidirectional Dimensionality Reduction
For computational and modeling convenience, the four-dimensional tensor is often reduced to some three-dimensional tensors from various directions ( shown in Figure 2). Below, we explain dimensionality reduction from three directions:
  • W/o User Dimension. For a user, his or her interactions in chronological order can form a sequence in each domain. These sequences reflect the user’s personalization as well as long and short-term preferences in different domains, and can further serve as the basis for constructing a local graph in subsequent work.
@@
  • W/o Domain Dimension. Blending two or more domains into a hybrid domain, a model can clarify the passing relationships between items originally in different domains. Many existing studies treat the hybrid domain as an independent domain, and design model fusion structures that incorporate all three domains in parallel, which is shown in Section 3.1.
@@ 2.3.1 Pure Sequential Representation
Arranging the interacted items of a user in chronological order yields a sequence. For each user $u$, we define $S^{A}_{u}=\{i^{A}_{1},i^{A}_{2},\dots\}$ and $S^{B}_{u}=\{i^{B}_{1},i^{B}_{2},\dots\}$ as his or her interaction sequences in domain A and domain B, respectively. If $u$ is an overlapping user who has interactions both in domain A and domain B, we can mix $S^{A}_{u}$ and $S^{B}_{u}$ to a hybrid sequence in chronological order (e.g., $S^{hybrid}_{u}=\{i^{A}_{1},i^{B}_{1},i^{A}_{2},i^{A}_{3},i^{B}_{2},\dots\}$), which is viewed as a sequence of the hybrid domain. From a pure sequential perspective, we can substitute the four-dimensional tensor in Eq.(1) with the set of sequences in each domain, as follows,
@@ 2.3.2 Graph-Encoded Sequential Representation
Some researchers Guo et al. (2021); Zheng et al. (2022); Cao et al. (2022); Zhang et al. (2023b) turn to construct directed graphs $G=\{V,E\}$ to model sequential information, where $V$ is a set of items that have been interacted with and $E$ is the edges that represent relations of the serial relationship from item to item. Constructing a user’s sequence within a domain or a sequence within a session as a local graph is a common practice Zheng et al. (2022); Chen et al. (2021). In the global graph construction scenario, $E$ is also utilized to denote the relations between users and items Xu et al. (2023c). We represent the raw data in its graph-encoded form to extend Eq.(1), where $G_{l}$ denotes all local graphs and $G_{g}$ denotes the global graph, as follows,
@@ 2.3.3 Side Information
More and more researchers consider incorporating side information to enrich the semantic representation of users’ historical behaviors. Here, we divide the most common side information into three categories: time, text (e.g., user/item context and review), and knowledge graph.
Time. Previously the timestamps are only used to order the items, but some researchers Xiao et al. (2023); Wang et al. (2022); Guo et al. (2022) define $t_{ij}=|t_{i}-t_{j}|$ to model the time interval and apply it to the subsequent learning of a model, where $t_{i}$ and $t_{j}$ are the timestamps that the item $i$ and the item $j$ are interacted with, respectively. Besides, we can also mine more information about time, such as periodicity and the duration of the interaction, etc.
@@
Knowledge Graph. Knowledge graphs Bi et al. (2020); Ma et al. (2022) deal with information from both entity and relationship views. It not only contains structural information between nodes but also implies some relationship descriptions. A knowledge graph defines an entity set $E^{KG}$ and a relation set $R^{KG}$, which consists of multiple entity-relation-entity triples $<e_{i},r,e_{j}>$ (e.g., $<e_{i_{1}^{A}},Is\_the\_same\_category,e_{i_{1}^{B}}>$ meaning that the entity $e_{i_{1}^{A}}$ from domain A has the same category as entity $e_{i_{1}^{B}}$ from domain B).
Considering the aforementioned side information, we can extend Eq.(1) as follows,
@@ 2.3.4 Pre-trained Features
Considering the privacy issues in two or more domains, some works Ding et al. (2023); Zhang et al. (2023a); Lei et al. (2021) use model-trained features as the input from the source domain, instead of raw data. Ding et al. (2023) directly combines the pre-trained user features from one domain with another domain. Zhang et al. (2023a) utilizes federated learning to fetch the global representation from the server, and then injects it into the local model.
@@ 3 Macro-View: What Structure Is Used to Fuse the CDSR Information?
In CDSR, the fusion structure serves as the primary skeleton of a model, providing pathways for separating and aggregating these features. In this section, we first describe multi-level fusion structures and then elaborate on the bridge for inter-domain fusion.
@@ 3.1.1 One-Level Fusion Structures
In earlier works, to combine the cross-domain information, there are a lot of works that first learn user preferences from domain A and domain B separately, and then fuse the representations of the two domains using various operations (shown in Figure 3(a)). These operations include, but are not limited to, concatenation Lei et al. (2021); Guo et al. (2023c), summation Alharbi and Caragea (2021); Ding et al. (2023), multi-layer perceptron (MLP) Bi et al. (2020), and some attention mechanisms Ouyang et al. (2020); Li et al. (2021).
Additionally, some works employ transfer learning to transfer knowledge from a source domain to a target domain Chen et al. (2021); Liu and Zhu (2021), or train a discriminator to bridge the representations of two domains with the idea of adversarial learning Li et al. (2022), shown in Figure 3(b).
@@ 3.1.2 Two/Three-Level Fusion Structures
Considering the potential for further exploration and advancement of fusion structures, some works combine the single domain and the hybrid domain with a multi-level fusion structure. As shown in Figure 3(e), in order to predict the next item in a separate domain, the hybrid domain is utilized as the main sharer Cao et al. (2022); Wang et al. (2022) or the bridge Zheng et al. (2022) to transfer knowledge from another domain. There are also some works Ye et al. (2023) that choose to combine domain A and domain B first and then fuse the hybrid information (i.e., Figure 3(f)).
@@ 3.1.3 Discussion
Although there are various multi-level fusion structures, it does not mean that a more complex structure will result in a better performance. The simple fusion structures, i.e., one-level fusion structures, are easy to implement but may not be conducive to comprehensive modeling of domain-specific and domain-generic features due to the limitation of the degree of fusion. While improving effectiveness, multi-level fusion structures bring increased complexity and reduced interpretability.
@@ 3.2 Bridges for Cross-Domain Fusion
When comes to cross-domain fusion, it is vital to clarify the bridges for information sharing between domains. We divide those bridges into three categories: same users, contexts, and clusters/groups, as shown in Figure 4. The first is prevalent in scenarios with overlapping users, while the latter two are applied in scenarios without overlapping users.
  • The Same Users. If some same users exist, then they are often the first choice as the bridge. Some works only focus on the fully overlapping users, neglecting non-overlapping users, while others Li et al. (2022) consider both cases.
  • Contexts. If there are no common users, it is feasible to leverage the semantic similarity of natural language. For instance, we can replace traditional item IDs with item contexts to explore similarities between items of different domains Liu et al. (2023).
@@
Different bridges exhibit distinct characteristics, each suitable for different scenarios. Works relying on overlapping users often perform well but their application scenarios may face limitations due to the sparse real-world data. Conversely, works built on non-overlapping users can be applied to a broader range of scenarios but may have limited performance.
@@ 4 Micro-View: What Technologies Are Used to Address the CDSR Problem?
In this section, we take a more tangible perspective to summarize the technologies adopted by the existing models in addressing the challenges. In conjunction with Table 2 and Figure 5, we elaborate on the basic technologies and the auxiliary learning technologies, respectively.
@@ 4.1.2 Attention Mechanisms in CDSR
Attention mechanisms can selectively focus on the importance of different parts and weight their contributions accordingly, enabling effective aggregation. Attention mechanisms are utilized in two main ways in CDSR modeling. One is to use an attention-based encoder (e.g., Transformer Vaswani et al. (2017), SASRec Kang and McAuley (2018) or multi-head attention blocks, etc.), replacing RNNs as a sequence encoder Ye et al. (2023); Xu et al. (2023c); Alharbi and Caragea (2022); Ding et al. (2023); Park et al. (2023). The other is to obtain learnable attention weights thus aggregating cross-domain information at multiple levels. For instance, MiNet Ouyang et al. (2020) designs item-level attention and interest-level attention to learn which items are more important and which of these items better matches a certain interest of the user. With representation-level attention, C2DSR Cao et al. (2022) fuses the sequential representation and the graph representation together. By adopting domain-level attention, Tri-CDR Ma et al. (2023) learns the weights of the features from different domains and then aggregates them. And MAN Lin et al. (2023a) also proposes a group-level attention to bridge the not-aligned information. In fact, attention strategies need to be designed from multiple levels which poses a challenge to researchers. However, it needs to use the interaction of all the positions when calculating the weights, so it contains a large number of parameters and may suffer from the issue of data sparsity.